The height of a right triangle is 8 feet more than twice the length of its base. Express the area of the triangle as a function of the length of its base, x, in feet.
Answer by Boreal(15235)

base=x
height is 2x+8 units are feet
Area is (1/2)bh=x(2x+8)(1/2)=x^2+4x sq ft